# Constraining color-charge effects of partonic energy loss with jet axis-based inclusive jet substructure measurement

### v1

@@ 3 Results and discussion
Figure 5 compares the resulting template fit curves (represented by green bands), referred to as the “two-component fit”, with the CMS $\Delta\text{j}$ distributions (black circles) for 0–10% centrality across various $p_{\text{T}}$ intervals. The ratios between the CMS data and simulation are displayed in the lower panels. The two-component fits capture the data trends at low $p_{\text{T}}$, the fit quality deteriorates in higher $p_{\text{T}}$ bins. It is worth noting that the discrepancies are most pronounced at high $\Delta\text{j}$ values, where little to no shape difference remains between the quark and gluon jet templates. These deviations suggest modifications beyond simple changes in the gluon fraction or simplistic $p_{\text{T}}$ shift are present in the data that cannot be captured by the vacuum pythia templates.
The figures also include unquenched pythia distributions (represented by the pink bands) and the modeling curve based on medium q/g fractions (shown as hatched lines). The unquenched pythia distributions exhibit the largest discrepancies with CMS data. The medium q/g modeling curve is constructed using pythia-based templates with a modified quark/gluon fractions, incorporating medium-induced modifications to jet functions CMS:2024koh ; mediumqg . This model describes the most central data points (for which predictions were available) across most of the measured $\Delta\text{j}$ range. However, at higher $\Delta\text{j}$ values in the higher $p_{\text{T}}$ bins, the model overestimates the width of the data distributions, even when considering predominantly narrow (quark-initiated) jets in simulation. This observation supports the conclusion that while the bulk of the distribution is well described, the large-$\Delta\text{j}$ behavior observed in CMS data cannot be explained solely by changes in the quark/gluon composition of the sample. Instead, this discrepancy highlights the need to account for medium-induced substructure modifications to more accurately describe the data.
Using the $p_{\text{T}}$ ranges obtained through minimum $\chi^{2}$ optimization, we recalculated the $\Delta\text{j}$ distributions in pythia and compared them with the CMS PbPb measurements for 0–10% centrality across four $p_{\text{T}}$ intervals (shown as colored bands in Fig. 5). Similar to the template fit curves, the $p_{\text{T}}$-shifted distributions align well with peripheral collision data within uncertainties but fail to fully reproduce the features of central collisions, particularly in the higher $p_{\text{T}}$ bins at larger $\Delta\text{j}$ values. Interestingly, the $p_{\text{T}}$-shifted distributions closely match predictions from the two-component fit and medium q/g models.
@@
Figure 5: The normalized $\Delta\text{j}$ distributions from pythia-based simulations using shifted inclusive $p_{\text{T}}$ and shifted quark/gluon $p_{\text{T}}$ (represented by color bands) are compared with CMS unfolded measurements in PbPb collisions at 0–10% centrality (shown as black circles) and with unquenched pythia simulations and a two-component fit (represented by pink and green color bands, respectively), as well as the medium q/g model (hatched lines), across various $p_{\text{T}}$ intervals. Only statistical uncertainties are shown for the simulation distributions. For the CMS data points, systematic uncertainties are indicated by rectangles (correlated uncertainties) and shaded areas (uncorrelated uncertainties), while vertical solid lines depict statistical uncertainties. In the lower panels, the color bands represent the total uncertainties in the data.
3.1 Limits on gluon fraction
@@
The left panel of Fig. 6 shows the extracted best-fit gluon fractions as a function of jet $p_{\text{T}}$, derived from the template fits to the CMS data, presented earlier (see Figs. 1 to 4). Since the template fitting attempts to capture all possible modifications by mere q/g fraction changes, the extracted gluon-like jet fraction values provide the lower bounds for what might be expected from realistic energy loss mechanisms. Any additional mechanisms contributing to the narrowing of the $\Delta\text{j}$ distribution (unrelated to changes in quark and gluon jet abundancies due to color-charge effects on energy loss) would likely increase this estimate.
Each variation of the $\Delta\text{j}$ distributions described in Section 5 of the CMS paper CMS:2024koh was used to estimate the systematic uncertainties associated with the extraction of gluon-like jet fractions. For each variation, the modified $\Delta\text{j}$ distribution was fitted, and the corresponding gluon-jet fraction was extracted. Deviations from the nominal value were treated as systematic uncertainties. When multiple variations were available for a given source, the largest deviation from the nominal was taken. The total uncertainty was calculated by combining the individual contributions in quadrature.
The right panel of Fig. 6 presents the ratio of the gluon fraction limits for each centrality interval relative to those in the 50–80% centrality range, emphasizing the centrality-dependent evolution of the observed variations. A lower limit on the gluon fractions is obtained for more central events, particularly at higher transverse momenta. This trend is more apparent in the ratio of the limits, where gluon-like jet fraction is reduced by approximately 50% for 0–10% central events. We find that the estimated gluon fraction from template fitting to the $\Delta\text{j}$ distributions is in agreement with the predictions from the medium q/g model for 0–10% central collisions CMS:2024koh ; mediumqg .
@@
3.2 Calculation of jet $R_{\text{AA}}$ using shifted $p_{\text{T}}$
To further explore the applicability of the $p_{\text{T}}$ shift estimates for jet quenching modeling, we construct a jet $R_{\text{AA}}$ estimate by dividing the jet counts in the shifted $p_{\text{T}}$ ranges by those in the corresponding nominal $p_{\text{T}}$ ranges in pythia. For the quark and gluon jet $p_{\text{T}}$ shifts, jets were first counted separately for each jet type, then combined, and the resulting counts were divided by those from the nominal $p_{\text{T}}$ ranges. The jet $R_{\text{AA}}$ estimate constructed using the inclusive jet $p_{\text{T}}$ shift is expected to provide a lower limit, while modeling the individual shifts to quark and gluon jet energies offers a more realistic estimate of energy loss. The resulting $R_{\text{AA}}$ estimates, shown in the left panel of Fig.7, are presented for both the inclusive jet $p_{\text{T}}$ shift and the independent quark/gluon jet $p_{\text{T}}$ shifts. The uncertainty on the $R_{\text{AA}}$ estimate is quantified by varying the $p_{\text{T}}$ shifts by minimizing $\chi^{2}+1$ and then calculating the difference between the resulting values. Remarkably, the constructed jet $R_{\text{AA}}$ estimates for both types of energy loss modeling align closely with previously measured jet $R_{\text{AA}}$ from the CMS and ATLAS experimentsCMS:2016uxf ; CMS:2021vui ; ATLAS:2018gwx . The agreement of the estimated quenching strength determined in this study, based on the jet axis decorrelation distributions, with earlier jet $R_{\text{AA}}$ measurements provides an independent validation of the approach used. The proposed studies of jet-axis-based substructure observables can thus offer additional constraints on medium-induced quark/gluon jet fraction modifications, leading to a more accurate assessment of energy loss in the medium.
@@ 3.2 Calculation of jet $R_{\text{AA}}$ using shifted $p_{\text{T}}$
3.3 Average fractional transverse momentum loss ($\Delta p_{\text{T}}\right>/p_{\text{T}})$
Alongside the the $R_{\text{AA}}$ estimate calculations, we also present the average fractional transverse momentum loss, ($\Delta p_{\text{T}}\right>/p_{\text{T}})$, separately for quark- and gluon-jets based on the $p_{\text{T}}$ shift values extracted for central 0–10% PbPb collisions, as summarized in Table 1. The uncertainty on these values was estimated in the same way as described earlier: minimizing $\chi^{2}+1$ and calculating the difference between the resulting values. The calculated $\Delta p_{\text{T}}\right>/p_{\text{T}}$ as a function of $p_{\text{T}}$ is shown in the right panel of Fig. 7. The results closely align with those reported in Ref. eg:partonloss , which used Bayesian analysis of multiple measurements across the studied $p_{\text{T}}$ ranges.
4 Summary
This paper presents a phenomenological study of jet axis decorrelation ($\Delta\text{j}$) between two types of jet axes for inclusive jets from lead-lead (PbPb) collisions at a center-of-mass energy of 5.02 TeV previously measured by the CMS experiment for different centrality selections and jet $p_{\text{T}}$ intervals. The fits to these distributions are performed using templates generated with the pythia8 event generator for gluon-initiated jets and jets from quarks of different flavors, attempting to understand the nature of the $\Delta\text{j}$ modifications observed.
@@ 4 Summary
It is found that the (unquenched) pythia8 predictions describe the peripheral CMS PbPb data but fail to capture the distributions from central events, due to progressive narrowing of $\Delta\text{j}$ distributions seen toward more central collisions. A simplified assumption about this narrowing as resulting entirely from changes in the quark/gluon jet fractions due to medium-induced energy loss was first tested. Limits on possible variations in jet sample composition were determined using pythia-based template fits to the CMS PbPb data. The fits result in a reduced fraction of gluon jets in the inclusive jet sample studied by CMS, with the most significant reduction observed for the highest-$p_{\text{T}}$ jets in the most central events. This observation is consistent with the expected higher in-medium energy loss by gluon-initiated jets compared to those originating from hard-scattered quarks.
Furthermore, possible differences in energy loss of quark vs. gluon jets were investigated by using $p_{\text{T}}$ shifts of the MC jet samples used for templates. Two different scenarios were probed, requiring a single $p_{\text{T}}$ shift for all (inclusive flavor) jets and shifting the quark- and gluon jets in pythia8 simulations independently. The study was performed for the most central CMS data, where medium-induced modification to the jet-axis decorrelation distributions is most significant. The results of this study yield, again, a higher energy loss for gluon than quark jets, consistent with theoretical expectations. The extracted $p_{\text{T}}$ shifts corresponding to the best description of the jet substructure data provide estimates of flavor-dependent energy loss as a function of jet $p_{\text{T}}$. To further test the validity of our approach and assumption, these $p_{\text{T}}$ shifts derived from CMS central PbPb collision data, the corresponding estimate for a jet nuclear modification factor $R_{\text{AA}}$ was calculated and shown to be consistent with the measured jet $R_{\text{AA}}$, providing a coherent jet quenching picture. Finally, the average fractional transverse momentum loss for jets initiated by different partons was calculated and compared with previous experimental and theoretical results.
@@ Acknowledgements.
The authors acknowledge Austin Baty, Hannah Bossi, Manuel Calderon De La Barca, Petar Maksimovic, Sevil Salur, and Leticia Cunqueiro Mendez for useful discussions and input during the development of this work. The authors thank Felix Ringer and Nobo Sato for providing the data points for Ref. mediumqg . This work is supported by the U.S. Department of Energy, grant No. DE-FG02-94ER40865.

### v2

@@ III Results and discussion
Figure 3 compares the resulting template fit curves (represented by green bands), referred to as the “two-component fit”, with the CMS $\Delta\text{j}$ distributions (black circles) for 0–10% centrality across various $p_{\text{T}}$ intervals. The ratios between the CMS data and simulation are displayed in the lower panels. The two-component fits capture the data trends at low $p_{\text{T}}$, but the fit quality deteriorates in higher $p_{\text{T}}$ bins. It is worth noting that the discrepancies are most pronounced at high $\Delta\text{j}$ values, where little to no shape difference remains between the quark and gluon jet templates. These deviations suggest modifications beyond simple changes in the gluon fraction or simplistic $p_{\text{T}}$ shifts are present in the data that cannot be captured by the vacuum pythia templates.
@@
Figure 3: The normalized $\Delta\text{j}$ distributions from pythia-based simulations using shifted inclusive $p_{\text{T}}$ (blue bands) and shifted quark/gluon $p_{\text{T}}$ (red bands) are compared with CMS unfolded measurements in PbPb collisions at 0–10% centrality (black circles). The results are also compared with unquenched pythia simulations (yellow dashed-line bands), a two-component fit (green bands), and the medium q/g model (hatched lines) across various $p_{\text{T}}$ intervals. Only statistical uncertainties are shown for the simulation distributions. For the CMS data points, systematic uncertainties are indicated by rectangles (correlated uncertainties) and shaded areas (uncorrelated uncertainties), while vertical solid lines depict statistical uncertainties. In the lower panels, the grey bands correspond to the total uncertainties in the data.
The figures also include unquenched pythia distributions (represented by the yellow dashed lines) and the modeling curve based on medium q/g fractions (shown as hatched lines). The unquenched pythia distributions exhibit the largest discrepancies with CMS data. The medium q/g modeling curve is constructed using pythia-based templates with modified quark/gluon fractions, incorporating medium-induced modifications to jet functions [CMS:2024koh, mediumqg]. This model describes the most central data points (for which predictions were available) across most of the measured $\Delta\text{j}$ range. However, at higher $\Delta\text{j}$ values in the higher $p_{\text{T}}$ bins, the model overestimates the width of the data distributions, even when considering predominantly narrow (quark-initiated) jets in simulation. This observation supports the conclusion that while the bulk of the distribution is well described, the large-$\Delta\text{j}$ behavior observed in CMS data cannot be explained solely by changes in the quark/gluon composition of the sample. Instead, this discrepancy highlights the need to account for medium-induced substructure modifications to more accurately describe the data.
Using the shifted $p_{\text{T}}$ ranges obtained through minimum $\chi^{2}$ optimization, we recalculated the $\Delta\text{j}$ distributions in pythia and compared them with the CMS PbPb measurements for 0–10% centrality across four $p_{\text{T}}$ intervals (shown as the blue and red bands in Fig. 3). Similar to the template fit curves, the $p_{\text{T}}$-shifted distributions align well with peripheral collision data within uncertainties but fail to fully reproduce the features of central collisions, particularly in the higher $p_{\text{T}}$ bins at larger $\Delta\text{j}$ values. Interestingly, the $p_{\text{T}}$-shifted distributions closely match predictions from the two-component fit and medium q/g models, with better agreement observed for the independent shift scenario.
III.1 Limits on gluon fraction
@@
The left panel of Fig. 4 shows the extracted best-fit gluon fractions as a function of jet $p_{\text{T}}$, derived from the template fits to the CMS data, presented earlier (see Fig. 2). Since the template fitting attempts to capture all possible modifications by mere q/g fraction changes, the extracted gluon-like jet fraction values provide the lower bounds for what might be expected from realistic energy loss mechanisms. Any additional mechanisms contributing to the narrowing of the $\Delta\text{j}$ distribution (unrelated to changes in quark and gluon jet abundances due to color-charge effects on energy loss) would likely increase this estimate.
@@
Each variation of the $\Delta\text{j}$ distributions described in Section 5 of the CMS paper [CMS:2024koh] was used to estimate the systematic uncertainties associated with the extraction of gluon-like jet fractions. For each variation, the modified $\Delta\text{j}$ distribution was fitted, and the corresponding gluon-jet fraction was extracted. Deviations from the nominal value were treated as systematic uncertainties. When multiple variations were available for a given source, the largest deviation from the nominal was taken. The total uncertainty was calculated by combining the individual contributions in quadrature.
The right panel of Fig. 4 presents the ratio of the gluon fraction limits for each centrality interval relative to those in the 50–80% centrality range, emphasizing the centrality-dependent evolution of the observed variations. A lower limit on the gluon fractions is obtained for more central events, particularly at higher transverse momenta. This trend is more apparent in the ratio of the limits, where gluon-like jet fraction is reduced by approximately 50% for 0–10% central events. We find that the estimated gluon fraction from template fitting to the $\Delta\text{j}$ distributions is in agreement with the predictions from the medium q/g model for 0–10% central collisions [CMS:2024koh, mediumqg].
III.2 Calculation of jet $R_{\text{AA}}$ using shifted $p_{\text{T}}$
To further explore the applicability of the $p_{\text{T}}$ shift estimates for jet quenching modeling, we construct a jet $R_{\text{AA}}$ estimate by dividing the jet counts in the shifted $p_{\text{T}}$ ranges by those in the corresponding nominal $p_{\text{T}}$ ranges in pythia. For the quark and gluon jet $p_{\text{T}}$ shifts, jets were first counted separately for each jet type, then combined, and the resulting counts were divided by those from the nominal $p_{\text{T}}$ ranges. The jet $R_{\text{AA}}$ estimate constructed using the inclusive jet $p_{\text{T}}$ shift is expected to provide a lower limit, while modeling the individual shifts to quark and gluon jet energies offers a more realistic estimate of energy loss. The resulting $R_{\text{AA}}$ estimates, shown in the left panel of Fig.5, are presented for both the inclusive jet $p_{\text{T}}$ shift and the independent quark/gluon jet $p_{\text{T}}$ shifts. The uncertainty on the $R_{\text{AA}}$ estimate is quantified by varying the $p_{\text{T}}$ shifts by minimizing $\chi^{2}+1$ and then calculating the difference between the resulting values. Remarkably, the constructed jet $R_{\text{AA}}$ estimates for both types of energy loss modeling align closely with previously measured jet $R_{\text{AA}}$ from the CMS and ATLAS experiments[CMS:2016uxf, CMS:2021vui, ATLAS:2018gwx]. The agreement of the estimated quenching strength determined in this study, based on the jet axis decorrelation distributions, with earlier jet $R_{\text{AA}}$ measurements provides support for the plausibility of the approach developed. The proposed studies of jet-axis-based substructure observables can thus offer additional constraints on medium-induced quark/gluon jet fraction modifications, leading to a more accurate assessment of energy loss in the medium.
@@ III.2 Calculation of jet $R_{\text{AA}}$ using shifted $p_{\text{T}}$
III.3 Average fractional transverse momentum loss $(\left<\Delta p_{\text{T}}\right>/p_{\text{T}})$
Alongside the $R_{\text{AA}}$ estimates, we also present the average fractional transverse momentum loss, $\langle\Delta p_{\text{T}}\rangle/p_{\text{T}}$, separately for quark and gluon jets, based on the $p_{\text{T}}$ shift values extracted for central 0–10% PbPb collisions, as summarized in Table 1. The uncertainties on these values were estimated in the same manner as described earlier: by minimizing $\chi^{2}+1$ and calculating the difference between the resulting values. The calculated $\langle\Delta p_{\text{T}}\rangle/p_{\text{T}}$ as a function of $p_{\text{T}}$ is shown in the right panel of Fig. 5. The results closely align with those reported in Ref. [eg:partonloss], which employed a Bayesian analysis of multiple measurements across the studied $p_{\text{T}}$ ranges.
IV Summary
This paper presents a phenomenological study of jet axis decorrelation ($\Delta\text{j}$) between two types of jet axes for inclusive jets from lead-lead (PbPb) collisions at a center-of-mass energy of 5.02 TeV previously measured by the CMS experiment for different centrality selections and jet transverse momentum ($p_{\text{T}}$) intervals. The fits to these distributions are performed using templates generated with the pythia8 event generator for gluon-initiated jets and jets from quarks of different flavors, attempting to understand the nature of the $\Delta\text{j}$ modifications observed.
@@ IV Summary
It is found that the (unquenched) pythia8 predictions describe the peripheral CMS PbPb data but fail to capture the distributions from central events, due to the progressive narrowing of the $\Delta\text{j}$ distributions observed toward more central collisions. Limits on possible variations in the jet sample composition were determined using pythia-based two-component template fits to the CMS PbPb data. The fits result in a reduced fraction of gluon jets in the inclusive jet sample studied by CMS, with the most significant reduction observed for the highest-$p_{\text{T}}$ jets in the most central events. This observation is consistent with the expected higher in-medium energy loss of gluon-initiated jets compared to those originating from hard-scattered quarks.
Furthermore, two possible scenarios of jet energy loss were investigated using transverse momentum shifted selections of jets in MC samples. First, an inclusive $p_{\text{T}}$ shift was applied to all jets (regardless of flavor), and then differences in energy loss between quarks and gluons were probed by independently shifting the momenta of quark- and gluon-initiated jets in pythia8 simulations. The study was performed for the most central CMS data, where medium-induced modification to the jet-axis decorrelation distributions is most significant. The results of the inclusive $p_{\text{T}}$-shift study do not provide the best description of the measured distributions, indicating a disfavoring of a common energy loss scenario. The results of the independent quark- and gluon-jet energy variations study indicate a higher energy loss for gluon jets compared to quark jets, consistent with theoretical expectations. The extracted $p_{\text{T}}$ shifts corresponding to the best description of the jet substructure data provide estimates of flavor-dependent energy loss as a function of jet $p_{\text{T}}$. The gluon jet fractions extracted from this approach are found to be reduced by amounts similar to those from the vacuum expectations as estimated from the two-component fit. To further assess the plausibility of our approach and assumptions, these $p_{\text{T}}$ shifts derived from CMS central PbPb collision data, the corresponding estimate for a jet nuclear modification factor $R_{\text{AA}}$ was calculated and shown to be consistent with the measured jet $R_{\text{AA}}$, providing a coherent jet quenching picture. Finally, the average fractional transverse momentum loss for jets initiated by different partons was calculated and compared with previous experimental and theoretical results.
@@ Acknowledgements.
The authors acknowledge Austin Baty, Hannah Bossi, Manuel Calderon De La Barca, Petar Maksimovic, Sevil Salur, and Leticia Cunqueiro Mendez for useful discussions and input during the development of this work. The authors thank Felix Ringer and Nobo Sato for providing the data points for Ref. [mediumqg]. This work is supported by the U.S. Department of Energy, grant No. DE-FG02-94ER40865.